# Size Generalization of Graph Neural Networks on Biological Data: Insights and Practices from the Spectral Perspective

### v4

@@ 5.1 Setup
5.2 (RQ1) Is cycle information helpful for size-generalization?
In this section, we aim to evaluate whether cycle information helps enhance the size generalizability of GNNs. We answer this question by evaluating three strategies: +SIA, +SSL, and +AugCyc. The two latter strategies incorporate cycle information through either a self-supervised auxiliary task or direct augmentation and are detailed in Appendix D.
As mentioned in Section 2.2, we evaluate the size generalizability of GNNs by training them on small graphs and testing their graph classification performance on large graphs. Better size generalizability translates into better performance on large graphs. Table 3 showcases the size generalizability results for our three proposed cycle-aware strategies, which we evaluate across five distinct datasets and six different backbone models. Notably, to better compare different strategies, the last column gives the average improvements compared with the original model evaluated across all datasets.
First, all of the cycle-aware strategies consistently lead to improvements in large test datasets without sacrificing the performance on small graphs. On average, these enhancements can reach up to 8.4%, affirming the effectiveness of cycle information in improving GNN size generalizability, as discussed in Section 3. Second, it is worth noting that cycle lengths provide more valuable information for enhancing size generalizability of GNNs. This is evident from the consistently better performance of the strategies +AugCyc and +SIA compared to the +SSL strategy on large test graphs, which solely predicts whether a node belongs to a cycle in the auxiliary task. Third, the simple attention-based strategy +SIA achieves the best overall performance improvements in all scenarios. On average, the +SIA strategy enhances both in-distribution and out-of-distribution generalization, while +SSL and +AugCyc excel particularly in out-of-distribution generalization. Additionally, +SIA achieves the highest average improvements on large test graphs. We attribute this to the challenges GNNs face in effectively learning cycle information, as shown in recent literature  (Chen et al., 2020). Furthermore, in Appendix F, we conduct an ablation study further demonstrating that the attention mechanism, without considering the cycle information, cannot improve the size generalizability of GNNs.
5.3 (RQ2) Comparison with Other Baselines
We now compare our best-performing strategy, +SIA, with other approaches and present their respective graph classification performances in Table 3. We find that +SIA consistently achieves the best performance compared with other baseline methods. While +RPGNN, +SSR, and +CIGA also enhance the size generalizability of GNNs, the improvements are less pronounced than those of +SIA. Additionally, it’s worth noting that +CIGAv1 show sensitivity to hyperparameters. While not explicitly designed for size generalizability, the expressive model SMP demonstrates strong performance on the BBBP dataset due to its cycle detection ability, validating our empirical insights. However, though SMP excels in cycle detection, overall, +SIA consistently outperforms SMP.
Table 4: Size generalizability evaluated with other baselines, following the same rule as in Table 3. +SIA consistently and significantly outperforms other strategies regarding size generalizability.
@@
6 Related Work
Literature on GNNs’ size generalization presents conflicting views. Empirical studies highlight GNNs’ size generalizability in physics simulation (Sanchez-Gonzalez et al., 2020) and algorithmic reasoning (Xu et al., 2021). Levie et al. (Levie et al., 2021) theoretically showed that spectral GNNs robustly transfer between graphs with varied sizes when discretizing the same underlying space. Meanwhile, some arguments suggest that GNNs may require additional assistance to achieve size generalizability. For instance, Yan et al. (Yan et al., 2020) and Velivckovic et al. (Veličković et al., 2020) found that neural networks effectively generalize to larger graphs than those used in training when attention weights are properly supervised. Conversely, some argue that GNN performance degrades with size shifts between training and test data, leading to the proposal of various models to mitigate this challenge. For instance, Yehudai et al. (Yehudai et al., 2021) argued that this performance degradation can be attributed to the changes in the local degree patterns. Knyazev et al. (Knyazev et al., 2019) found that using attention with proper thresholding can improve the size generalizability of GNNs. Buffelli et al. (Buffelli et al., 2022) simulated a size shift in the training graphs via graph coarsening and proposed a regularization that makes the model robust to the shift. Bevilacqua et al. (Bevilacqua et al., 2021) used a causal model to learn approximately invariant representations that better extrapolate between train and test data. Chen et al. (Chen et al., 2022) utilized structural causal models for robust out-of-distribution generalization in graph data through invariant subgraph identification and label prediction. Chu et al. (Chu et al., 2023) proposed a Wasserstein barycenter matching (WBM) layer to address the slow uncontrollable convergence rate w.r.t. graph size. Zhou et al. (Zhou et al., 2022) studied the size OOD problem in the task of link prediction. Ji et al. (Ji et al., 2022) curated OOD datasets for AI-aided drug discovery. Our study stands out as the first to utilize spectral analysis to characterize the types of size-induced distribution shifts, shedding light on the underlying causes that hinder GNNs from effectively generalizing to large graphs. Some expressive models also exhibit robustness in size generalization. Murphy et al. (Murphy et al., 2019) proposed an expressive model-agnostic framework that learns graph representations invariant to graph isomorphism given variable-length inputs. Clement et al. (Vignac et al., 2020) proposed an expressive graph neural network that performs well on difficult structural tasks, such as cycle detection and diameter computation. Our study validates that expressive models excelling in cycle-related tasks demonstrate good size generalizability.
7 Conclusion
@@
In conclusion, our work extensively characterizes size-induced distribution shifts and evaluates their impact on GNNs’ generalizability to significantly larger test graphs compared ot the training set. Spectral analysis on real-world biological data reveals a strong correlation between graph spectrum and size, which hinders GNNs’ size generalization. We identify the pivotal role of cycle-related information in reducing spectral differences between small and alrge graphs. Motivated by these findings, we introduce a size-insensitive attention strategy that can be integrated with any GNN architectures. Our empirical findings suggest that cycle information is cruical for size generalization and our proposed method significantly enhances the size generalizability of GNNs. This research provides valuable insights for enhacing GNN generalization across varying graph sizes.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. Our research reveals important insights in the scientific domain, i.e., biology, contributing to the enhancement and progress of the field.
@@ Appendix A Proposition
In this Appendix, we provide theoretical reasoning for the proposition in Section 3.1.
Without loss of generality, the output of a spectral GNN at the $(l+1)$-th layer is given by: ${\mathbf{X}}^{(l+1)}=\sigma({\mathbf{U}}f({\mathbf{\Lambda}}){\mathbf{U}}^{T}{\mathbf{X}}^{(l)}{\mathbf{W}}^{(l)})$ (Section 2). To ease our analysis, we rewrite ${\mathbf{U}}$, ${\mathbf{\Lambda}}$ and ${\mathbf{X}}^{(l)}$ as: ${\mathbf{U}}=[{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}]$, ${\mathbf{\Lambda}}=\texttt{Diag}([\lambda_{1},\lambda_{2},\cdots,\lambda_{N}])$, and ${\mathbf{X}}^{(l)}=[{\mathbf{X}}^{(l)}_{1},{\mathbf{X}}^{(l)}_{2},\cdots,{\mathbf{X}}^{(l)}_{D}]$, respectively. ${\mathbf{U}}_{i}$ is the $i$-th column vector of ${\mathbf{U}}$, $\lambda_{i}$ is the $i$-th largest eigenvalue, $\texttt{Diag}(\cdot)$ creates a matrix with the vector as its diagonal elements, ${\mathbf{X}}^{(l)}_{i}$ is the $i$-th column vector of ${\mathbf{X}}^{(l)}$, and $D$ is the feature dimension at the $l$-th layer.
$\begin{split}{\mathbf{X}}^{(l+1)}&=\sigma([{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}]\cdot\texttt{Diag}([f(\lambda_{1}),f(\lambda_{2}),\cdots,f(\lambda_{N})])\cdot([{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}])^{T}\\ &\cdot[{\mathbf{X}}^{(l)}_{1},{\mathbf{X}}^{(l)}_{2},\cdots,{\mathbf{X}}^{(l)}_{D}]\cdot{\mathbf{W}}^{(l)})\\ &=\sigma([f(\lambda_{1}){\mathbf{U}}_{1},f(\lambda_{2}){\mathbf{U}}_{2},\cdots,f(\lambda_{N}){\mathbf{U}}_{N}]\cdot([{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}])^{T}\\ &\cdot[{\mathbf{X}}^{(l)}_{1},{\mathbf{X}}^{(l)}_{2},\cdots,{\mathbf{X}}^{(l)}_{D}]\cdot{\mathbf{W}}^{(l)}).\end{split}$ (1)
Since $\{{\mathbf{U}}_{i}:i=1,\cdots,N\}$ are the orthonormal basis of ${\mathbb{R}}^{N}$, ${\mathbf{X}}^{(l)}_{i}$ can be expressed as a linear combination of $\{{\mathbf{U}}_{i}\}$. Thus, suppose ${\mathbf{X}}^{(l)}_{i}=\sum_{j=1,2,\cdots,N}{\alpha_{j}^{i}{\mathbf{U}}_{j}}$, then Equation 1 can be rewritten as:
$\begin{split}{\mathbf{X}}^{(l+1)}&=\sigma([f(\lambda_{1}){\mathbf{U}}_{1},f(\lambda_{2}){\mathbf{U}}_{2},\cdots,f(\lambda_{N}){\mathbf{U}}_{N}]\cdot([{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}])^{T}\\ &\cdot[\sum_{j=1,2,\cdots,N}{\alpha_{j}^{1}{\mathbf{U}}_{j}},\cdots,\sum_{j=1,2,\cdots,N}{\alpha_{j}^{i}{\mathbf{U}}_{j}},\cdots]\cdot{\mathbf{W}}^{(l)})\\ &=\sigma\left([f(\lambda_{1}){\mathbf{U}}_{1},f(\lambda_{2}){\mathbf{U}}_{2},\cdots,f(\lambda_{N}){\mathbf{U}}_{N}]\cdot\begin{bmatrix}\alpha_{1}^{1}&\alpha_{1}^{2}&\cdots&\alpha_{1}^{D}\\ \alpha_{2}^{1}&\alpha_{2}^{2}&\cdots&\alpha_{2}^{D}\\ \cdots&\cdots&\cdots&\cdots\\ \alpha_{N}^{1}&\alpha_{N}^{2}&\cdots&\alpha_{N}^{D}\end{bmatrix}\cdot{\mathbf{W}}^{(l)}\right)\\ &=\sigma([\sum_{j=1,\cdots,N}{f(\lambda_{j})\alpha_{j}^{1}{\mathbf{U}}_{j}},\sum_{j=1,\cdots,N}{f(\lambda_{j})\alpha_{j}^{2}{\mathbf{U}}_{j}},\cdots,\sum_{j=1,\cdots,N}{f(\lambda_{j})\alpha_{j}^{D}{\mathbf{U}}_{j}}]\cdot{\mathbf{W}}^{(l)}).\end{split}$ (2)
@@
$\begin{split}{\mathbf{X}}^{(l+1)}&=\sigma([\sum_{j=1,\cdots,N}{f(\lambda_{j})\texttt{COSINE}({\mathbf{U}}_{j},{\mathbf{X}}^{(l)}_{1})\cdot\frac{\lVert{\mathbf{X}}^{(l)}_{1}\rVert_{2}}{\sqrt{N}}\cdot(\sqrt{N}\cdot{\mathbf{U}}_{j})},\\ &\cdots,\sum_{j=1,\cdots,N}{f(\lambda_{j})\texttt{COSINE}({\mathbf{U}}_{j},{\mathbf{X}}^{(l)}_{D})\cdot\frac{\lVert{\mathbf{X}}^{(l)}_{D}\rVert_{2}}{\sqrt{N}}\cdot(\sqrt{N}\cdot{\mathbf{U}}_{j})}]\cdot{\mathbf{W}}^{(l)}).\end{split}$ (3)
The final graph representation is obtained through a global pooling (Section 2), where a set function is applied to each feature dimension. To ensure that the graph representation is irrelevant to the graph size, it is crucial for each column of ${\mathbf{X}}^{(l+1)}$ to be unaffected by the size. If the distributions of eigenvalues ($\lambda_{j}$), scaled eigenvectors ($\sqrt{N}\cdot{\mathbf{U}}_{j}$), and the scalar $\frac{\lVert{\mathbf{X}}^{(l)}_{i}\rVert_{2}}{\sqrt{N}}$, are uncorrelated with the graph size, then the graph representation will be size-invariant. It is worth noting that ${\mathbf{X}}^{(l)}_{i}$ contains $N$ elements and its 2-norm scales with $\sqrt{N}$; and $\lVert{\mathbf{U}}_{j}\rVert_{2}=1$ and its elements diminish with $\sqrt{N}$. Hence, we account for the scaling factor $\sqrt{N}$ in both terms. However, the distributions of eigenvalues (Section 3.2) and scaled eigenvectors in real biological graphs are influenced by the graph size. This poses a challenge for the size generalization of GNNs. Importantly, the difference in eigenvalue distributions cannot be attributed to the value range (Berlingerio et al., 2012), as the eigenvalues of a normalized Laplacian matrix and adjacency matrix are constrained within the intervals [0,2] and [-1,1] respectively.

### v5

@@ 5.1. Experimental Setup
Table 4. Size generalizability evaluated with other baselines, following the same rule as in Table 3. +SIA consistently and significantly outperforms other strategies regarding size generalizability.
@@
5.2. (RQ1) Is Cycle Information Helpful?
In this section, we aim to evaluate whether cycle information in our three proposed strategies (+SIA, +SSL, and +AugCyc) helps enhance the size generalizability of GNNs.
Setup. We evaluate the size generalizability of GNNs by training them on small graphs and testing their graph classification performance on large graphs. Better size generalizability translates into better performance on large graphs. Table 3 showcases the size generalizability results for our three proposed cycle-aware strategies, which we evaluate across five distinct datasets and six different backbone models. Notably, to better compare different strategies, the last column gives the average improvements compared with the original model evaluated across all datasets.
Results. First, all of the cycle-aware strategies consistently lead to improvements in large test datasets without sacrificing the performance on small graphs. On average, these enhancements can reach up to 8.4%, affirming the effectiveness of cycle information in improving GNN size generalizability, as discussed in Section 3. Second, it is worth noting that cycle lengths provide more valuable information for enhancing size generalizability of GNNs. This is evident from the consistently better performance of the strategies +AugCyc and +SIA compared to the +SSL strategy on large test graphs, which solely predicts whether a node belongs to a cycle in the auxiliary task. Third, the simple attention-based strategy, +SIA, achieves the best overall performance improvements in all scenarios. On average, the +SIA strategy enhances both in-distribution and out-of-distribution generalization, while +SSL and +AugCyc excel particularly in out-of-distribution generalization. Additionally, +SIA achieves the highest average improvements on large test graphs. We attribute this to the challenges GNNs face in effectively learning cycle information, as shown in recent literature  (Chen et al., 2020). Also, in Appendix F, we conduct an ablation study further demonstrating that the attention mechanism, without considering the cycle information, cannot improve the GNN size generalizability.
Discussion. The three strategies examined in this study primarily highlight the importance of cycle information in achieving size generalization. However, we do not recommend the combined use of all three strategies. Table 3 demonstrates that these strategies differ in their effectiveness for in-distribution generalization, with +AugCyc sometimes leading to performance degradation. Notably, +SSL can be considered a subset of +SIA, as +SIA incorporates more comprehensive information, including both cycle length and is_cycle, into its mechanism. Therefore, we recommend using +SIA in practice due to its broader scope, while the other two strategies serve to illustrate the practical value of cycle information.
Table 5. Performance of baseline and +SIA across various model architectures in the synthetic dataset.
Method MLP-small MLP-large GCN-small GCN-large GAT-small GAT-large GIN-small GIN-large FAGCN-small FAGCN-large Avg-small Avg-large
Original 63.61±4.50 59.72±8.50 64.75±1.56 32.61±8.36 65.48±2.17 55.14±5.64 67.01±1.60 44.16±6.12 65.32±1.86 61.45±1.88 65.23 50.62
SIA 63.82±4.30 62.56±8.48 72.88±3.67 69.61±9.38 64.52±4.56 63.72±6.33 74.13±4.23 67.56±8.67 71.80±3.42 69.07±3.11 69.43 66.50
5.3. (RQ2) Comparison with Other Baselines
We now compare our best-performing strategy, +SIA, with other approaches and present their respective graph classification performances in Table 4. Averaged improvements are presented in Table 11 in Appendix E. We find that +SIA consistently achieves the best performance compared with other baseline methods. While +RPGNN, +SSR, and +CIGA also enhance the size generalizability of GNNs, the improvements are less pronounced than those of +SIA. Additionally, it’s worth noting that +CIGAv1 shows sensitivity to hyperparameters. While not explicitly designed for size generalizability, the expressive model SMP demonstrates strong performance on the BBBP dataset due to its cycle detection ability, validating our empirical insights. However, though SMP excels in cycle detection, overall, +SIA consistently outperforms SMP. In addition, our proposed strategies result in minimal increases in training time compared to the baselines. We present the per-epoch runtime statistics in Table 10 and the average results in Table 11.
5.4. (RQ3) Effectiveness on Large Dataset
We evaluate the effectiveness of our method (+SIA) on the HIV dataset from Hu et al. (2020). Using the ROC-AUC metric (Hu et al., 2020; Chen et al., 2022), we find that +SIA consistently improves performance across all backbone models, as shown in Table 6. On average, +SIA yields a +0.15 improvement on small test sets and a +3.63 improvement on large test sets, demonstrating its efficacy on real benchmarks.
Table 6. Performance comparison of different models with and without SIA on large dataset. -S and -L denote small and large test set performance, respectively.
Models MLP GCN GAT GIN FAGCN GNNML3
Original-S 67.95±0.08 71.29±0.35 73.34±1.17 71.04±0.34 68.35±0.50 74.22±0.61
Original-L 38.97±1.07 38.35±1.29 41.22±1.26 43.55±0.44 50.92±0.33 47.44±0.84
+SIA-S 70.64±1.14 70.77±1.42 72.33±0.32 69.62±0.43 69.53±0.39 74.52±1.55
+SIA-L 40.93±0.79 43.24±1.44 46.72±1.05 43.55±0.45 52.58±1.45 54.68±0.44
5.5. (RQ4) Verification on Synthetic Dataset
Going beyond real datasets, we further evaluate various models on synthetic data (Section 5.1) and verify our findings in controlled settings. The results, shown in Table 5, highlight two key observations. First, the original model struggles to generalize across different graph sizes. Second, the effectiveness of +SIA is even more pronounced in this controlled synthetic setting.
Additionally, we extend the aligning/breaking cycle analysis from Table 2 to the synthetic dataset. As shown in Table 7, while the absolute spectral differences between graphs of different and similar sizes remain largely unchanged, the relative spectral differences exhibit clear variations. These observations reinforce our claims.
Table 7. Effects of breaking and aligning cycles on synthetic datasets, evaluated using the same metrics as in Table 2.
Original Breaking Cycles Aligning Cycles
Different Size 0.0051 0.0049 0.0050
Similar Size 0.0016 0.0011 0.0021
Relative Difference 220% 356% (+136%) 134% (-86%)
6. Related Work
Literature on GNNs’ size generalization presents conflicting views. On one hand, empirical studies highlight GNNs’ size generalizability in physics simulation (Sanchez-Gonzalez et al., 2020) and algorithmic reasoning (Xu et al., 2021). Levie et al. (Levie et al., 2021) theoretically showed that spectral GNNs robustly transfer between graphs with varied sizes when discretizing the same underlying space.
On the other hand, some works suggest that GNNs may require additional information to achieve size generalizability. For instance, Yan et al. (Yan et al., 2020) and Velivckovic et al. (Veličković et al., 2020) found that neural networks effectively generalize to larger graphs than those used in training when attention weights are properly supervised. In contrast, Yehudai et al. (Yehudai et al., 2021) argued that performance degradation is linked to changes in degree patterns within multi-hop neighborhoods, known as ‘d-patterns’. However, our findings indicate that d-patterns do not adequately capture variations in cycle statistics, a critical factor for generalizing GNNs to larger biological graphs (Appendix K). Additionally, our analysis shows that degree distribution does not correlate distinctly with graph size in biological datasets (Appendix J).
In general, various methodologies have been proposed to address performance degradation from size shifts, including attention with proper thresholding (Knyazev et al., 2019), stochastic anchoring (Trivedi et al., 2024), regularization that improves model robustness to size shifts via graph coarsening (Buffelli et al., 2022), a causal model to learn invariant representations for approximately better extrapolation between training and test data (Bevilacqua et al., 2021), structural causal models for robust out-of-distribution generalization through invariant subgraph identification (Chen et al., 2022), a Wasserstein barycenter matching (WBM) layer to improve convergence rates with respect to graph size (Chu et al., 2023), and a disentanglement learning framework to encourage size-invariant learning (Huang et al., 2024). Our study stands out as the first to utilize spectral analysis to characterize the types of size-induced distribution shifts, shedding light on the underlying causes that hinder GNNs from effectively generalizing to large graphs.
Some expressive models also show promise in addressing size generalization. For example, Murphy et al. (Murphy et al., 2019) proposed a model-agnostic framework that learns graph representations invariant to graph isomorphism given variable-length inputs, while Clement et al. (Vignac et al., 2020) proposed an expressive GNN that performs well on difficult structural tasks, such as cycle detection and diameter computation. Our study validates that expressive models excelling in cycle-related tasks demonstrate good size generalizability. Other related works include studying the size OOD problem in the task of link prediction (Zhou et al., 2022), curating OOD datasets for AI-aided drug discovery (Ji et al., 2023), and improving uncertainty estimates in GNNs under various distribution shifts, including shifts in graph size (Trivedi et al., 2024).
7. Conclusion
@@
Our work extensively characterized size-induced distribution shifts and evaluated their impact on GNNs’ generalizability to significantly larger test graphs compared to the training set. Through spectral analysis on real-world biological data, we uncovered a strong correlation between graph spectrum and size, which hinders GNNs’ size generalization. We further identified the pivotal role of cycle-related information in reducing spectral differences between small and large graphs. Motivated by these findings, we introduced three cycle-aware strategies that can be integrated with any GNN architectures. Our empirical findings suggest that cycle information is crucial for size generalization and our proposed method significantly enhances the size generalizability of GNNs. Overall, this work provides practical insights for enhacing GNN generalization across varying graph sizes. An interesting future direction is to extend this work to domains beyond biology.
Acknowledgements
We thank the anonymous reviewers for their constructive feedback. This material is based upon work supported by the National Science Foundation under IIS 2212143, and CAREER Grant No. IIS 1845491. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation or other funding parties.
@@ Appendix A Proposition
In this section, we provide theoretical reasoning for the proposition in Section 3.3.
[…]
$\begin{split}{\mathbf{X}}^{(l+1)}&=\sigma([{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}]\cdot\texttt{Diag}([f(\lambda_{1}),f(\lambda_{2}),\cdots,f(\lambda_{N})])\\ &\cdot[{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}]^{T}\cdot[{\mathbf{X}}^{(l)}_{1},{\mathbf{X}}^{(l)}_{2},\cdots,{\mathbf{X}}^{(l)}_{D}]\cdot{\mathbf{W}}^{(l)})\\ &=\sigma([f(\lambda_{1}){\mathbf{U}}_{1},f(\lambda_{2}){\mathbf{U}}_{2},\cdots,f(\lambda_{N}){\mathbf{U}}_{N}]\cdot([{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}])^{T}\\ &\cdot[{\mathbf{X}}^{(l)}_{1},{\mathbf{X}}^{(l)}_{2},\cdots,{\mathbf{X}}^{(l)}_{D}]\cdot{\mathbf{W}}^{(l)}).\end{split}$
Since $\{{\mathbf{U}}_{i}:i=1,\cdots,N\}$ form an orthonormal basis of ${\mathbb{R}}^{N}$, ${\mathbf{X}}^{(l)}_{i}$ can be expressed as a linear combination of $\{{\mathbf{U}}_{i}\}$. Thus, supposing that ${\mathbf{X}}^{(l)}_{i}=\sum_{j=1}^{N}{\alpha_{j}^{i}{\mathbf{U}}_{j}}$, Equation 8 can be rewritten as:
$\begin{split}{\mathbf{X}}^{(l+1)}&=\sigma([f(\lambda_{1}){\mathbf{U}}_{1},f(\lambda_{2}){\mathbf{U}}_{2},\cdots,f(\lambda_{N}){\mathbf{U}}_{N}]\cdot([{\mathbf{U}}_{1},{\mathbf{U}}_{2},\cdots,{\mathbf{U}}_{N}])^{T}\\ &\cdot[\sum_{j=1}^{N}{\alpha_{j}^{1}{\mathbf{U}}_{j}},\cdots,\sum_{j=1}^{N}{\alpha_{j}^{i}{\mathbf{U}}_{j}},\cdots]\cdot{\mathbf{W}}^{(l)})\\ &=\sigma\left([f(\lambda_{1}){\mathbf{U}}_{1},f(\lambda_{2}){\mathbf{U}}_{2},\cdots,f(\lambda_{N}){\mathbf{U}}_{N}]\cdot\begin{bmatrix}\alpha_{1}^{1}&\alpha_{1}^{2}&\cdots&\alpha_{1}^{D}\\ \alpha_{2}^{1}&\alpha_{2}^{2}&\cdots&\alpha_{2}^{D}\\ \cdots&\cdots&\cdots&\cdots\\ \alpha_{N}^{1}&\alpha_{N}^{2}&\cdots&\alpha_{N}^{D}\end{bmatrix}\cdot{\mathbf{W}}^{(l)}\right)\\ &=\sigma([\sum_{j=1}^{N}{f(\lambda_{j})\alpha_{j}^{1}{\mathbf{U}}_{j}},\sum_{j=1}^{N}{f(\lambda_{j})\alpha_{j}^{2}{\mathbf{U}}_{j}},\cdots,\sum_{j=1}^{N}{f(\lambda_{j})\alpha_{j}^{D}{\mathbf{U}}_{j}}]\cdot{\mathbf{W}}^{(l)}).\end{split}$
@@
$\begin{split}{\mathbf{X}}^{(l+1)}&=\sigma([\sum_{j=1,\cdots,N}{f(\lambda_{j})\texttt{COSINE}({\mathbf{U}}_{j},{\mathbf{X}}^{(l)}_{1})\cdot\frac{\lVert{\mathbf{X}}^{(l)}_{1}\rVert_{2}}{\sqrt{N}}\cdot(\sqrt{N}\cdot{\mathbf{U}}_{j})},\\ &\cdots,\sum_{j=1,\cdots,N}{f(\lambda_{j})\texttt{COSINE}({\mathbf{U}}_{j},{\mathbf{X}}^{(l)}_{D})\cdot\frac{\lVert{\mathbf{X}}^{(l)}_{D}\rVert_{2}}{\sqrt{N}}\cdot(\sqrt{N}\cdot{\mathbf{U}}_{j})}]\cdot{\mathbf{W}}^{(l)}).\end{split}$
Here, $\frac{\lVert{\mathbf{X}}^{(l)}_{i}\rVert_{2}}{\sqrt{N}}=C$ holds due to feature normalization (Singh and Singh, 2022).